OpenStudy (anonymous):

Giving medal and best answer for help! The solution of 3/(n squared) - 1/(n) = -1/(2n) is n = -2. True False

OpenStudy (zzr0ck3r):

3/n^2=-1/(2n) sub in -2 for n 3/4 = -1/(-4) = 1/4 3/4=1/4 is not true so false

OpenStudy (anonymous):

thank you! @zzr0ck3r

OpenStudy (zzr0ck3r):

wait wait wait

OpenStudy (zzr0ck3r):

I swear I know what im doing and that is the last time:)

OpenStudy (zzr0ck3r):

wt heck

OpenStudy (zzr0ck3r):

I didn't mean to delete it

OpenStudy (zzr0ck3r):

it fase

OpenStudy (zzr0ck3r):

false

OpenStudy (anonymous):

haha no problem at all! Thanks for the help!